OpenStudy (korosh23):

***Physics 12 Question*** Two planets A and B have the same mass. However, the gravitational field strength on the surface of the planet A is 1.20 times the gravitational field strength on the surface of planet B. How does the radius of planet A compare with the radius of planet B? Please explain and show me your work. Thank you.

OpenStudy (michele_laino):

the gravity of planet A, is: \[\Large {g_A} = G\frac{{{M_A}}}{{R_A^2}}\] where the gravity on the Surface of planet B, is: \[\Large {g_B} = G\frac{{{M_B}}}{{R_B^2}}\] now, if we divide such equation side by side together, we get: \[\Large \frac{{{g_A}}}{{{g_B}}} = \frac{{{M_A}}}{{R_A^2}} \cdot \frac{{R_B^2}}{{{M_B}}} = {\left( {\frac{{{R_B}}}{{{R_A}}}} \right)^2}\] since \(M_A=M_B\) now, we use the condition \(g_A=1.2\;g_B\), so we get: \[\Large\frac{6}{5} = {\left( {\frac{{{R_B}}}{{{R_A}}}} \right)^2}\] so, what can you conclude?